A worker on a roof, 20.8 m above the ground, wants to drop a brick onto the back of a moving truck. The
truck is moving at a constant velocity of 5.0 m/s alongside the building. The back of the truck is 1.2 m
above the ground. How far from the point of impact must the truck be, the instant the brick is dropped, so
that the brick will land on the back of the truck?
1
Expert's answer
2019-04-15T09:47:12-0400

Time of flight:


"t= \sqrt{\frac{2(H-h)}{g}}= \sqrt{\frac{2(20.8-1.2)}{9.8}}=2 s"

Needed distance is


"s=vt=(5)(2)=10m."